OpenStudy (anonymous):

Find the probabilities of getting three 3's and then a 4 or a 5 in four rolls of a balanced die.

OpenStudy (anonymous):

Getting a three, four and five out of a balanced die is \[\frac{3}{6},\frac{4}{6},\frac{5}{6}\] respectfully

OpenStudy (anonymous):

no that is not right

OpenStudy (anonymous):

rolling 1 or 4 or 5 or any number has the same chance. What is it? What is the chance of rolling 3?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

Sorry I mean 1/6

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

So in this case it is saying that you roll 4 times. And what is the chance that you will get 3,3,3,4 or 3,3,3,5

OpenStudy (anonymous):

These are called independent events. That means that the first rolls result does not effect the second or third or fourth rolls result.

OpenStudy (anonymous):

Probability of rolling a 3 in a side sided die is 1/6. Probabilty of rolling three 3s in a row is \[\frac{1}{6}*\frac{1}{6}*\frac{1}{6}=\frac{1}{6^3}\] Probabilty of rolling a 4 or a 5 is the combination of the probability of rolling a 4 plus the probability of rolling a 5 \[\frac{1}{6}+\frac{1}{6}=\frac{2}{6}=\frac{1}{3}\] The probability of rolling a 4 or a 5 after you roll three 3s is \[\frac{1}{6}*\frac{1}{6}*\frac{1}{6}*\frac{1}{3}=\frac{1}{3*6^3}\]

OpenStudy (anonymous):

If two or more events are independent than the probability of both happening at the same time is the multiple of each probability.

OpenStudy (anonymous):

Is it clear what 2le wrote? It is correct

OpenStudy (anonymous):

Yes, I was able to get up to the second part of his answer, and couldn't figure out the last part but now I understand it. Thank you both.